davif:

The graph below shows the prices of different numbers of swimsuits at a store: A graph is shown. The values on the x axis are 0, 2, 4, 6, 8, and 10. The values on the y axis are 0, 23, 46, 69, 92, and 115. Points are shown on ordered pairs 0, 0 and 2, 23 and 4, 46 and 6, 69 and 8, 92. These points are connected by a line. The label on the x axis is Number of Swimsuits. The title on the y axis is Price in dollars. Which equation can be used to determine p, the cost of s swimsuits? p = 11.50 + s s = 11.50 + p s = 11.50p p = 11.50s

davif:

help

jhonyy9:

so x is the number of swimsuits y is the price in dollars x values 0,2,4,6,8,10 y values 0,23,46,69,92,115 Which equation can be used to determine p ? p = 11.50 + s s = 11.50 + p s = 11.50p p = 11.50s so look pls how you think in case of what equation we will can get using the given x values what is s inside equations the necessary y values from these given dates what mean p ?

jhonyy9:

i think the easyst will be to calcule and look the last one how you get from 11,50 *s the 115 how many need being s ?

jhonyy9:

@Laylalyssa do you agree this ?

jhonyy9:

ty

Laylalyssa:

yes..I think so..

darkknight:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Laylalyssa yes..I think so.. \(\color{#0cbb34}{\text{End of Quote}}\) well if someone tags you for an opinion you should state y next time. x values 0,2,4,6,8,10 y values 0,23,46,69,92,115 now @jhonyy9 we see that 0,0 is a point so no y-intercept, we are relating the price of swimsuits, (11.5 each) so I agree that the last one is correct

Laylalyssa:

mb😫

darkknight:

its perfectly fine, just against QC rules yk to not repeat or contribute unless clarifying something. @jhonyy9 i don't understand y u tag laylalyssa and other people? also can this be a dm matter so that the post isn't spammed?

Laylalyssa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight its perfectly fine, just against QC rules yk to not repeat or contribute unless clarifying something. @jhonyy9 i don't understand y u tag laylalyssa and other people? also can this be a dm matter so that the post isn't spammed? \(\color{#0cbb34}{\text{End of Quote}}\) why can't he tag laylalyssa and other ppl?

supie:

@darkknight

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight ? \(\color{#0cbb34}{\text{End of Quote}}\) wuts ur problem dude? everything jhonyy9 doesn you have a problem with it like sit down breathe relax

Laylalyssa:

exactly

darkknight:

all i said is that if you can't really contribute don't reply, just to help the spam, i have no problem and stop commenting off topic/spam. ill delete mine kindly delete yours : )

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight all i said is that if you can't really contribute don't reply, just to help the spam, i have no problem and stop commenting off topic/spam. ill delete mine kindly delete yours : ) \(\color{#0cbb34}{\text{End of Quote}}\) uhh no you said more than that and also he asked if she agreed she said yes i agree what is the problem with that?

supie:

@darkknight -_-

darkknight:

its a rule now lets chill, u ain't having a good iamge for urself

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight its a rule now lets chill, u ain't having a good iamge for urself \(\color{#0cbb34}{\text{End of Quote}}\) its not a rule that you cant ask someone for their opinion and they cant give their opinion

supie:

"i don't understand y u tag laylalyssa and other people?" why cant jhonyy9 tag people?

darkknight:

@dude @Vocaloid @shadow I don't remember which but one of you said that but there was a rule about inadequate responses. Would like you to clarify to them about that. Jhonny can tag i never said he can't. just don't "spam"

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight @dude @Vocaloid @shadow I don't remember which but one of you said that but there was a rule about inadequate responses. Would like you to clarify to them about that. Jhonny can tag i never said he can't. just don't "spam" \(\color{#0cbb34}{\text{End of Quote}}\) He asked for her opinion she was just confirming that it's right, which was her opinion, that it was right i dont understand what your problem is with that?

darkknight:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie \(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight @dude @Vocaloid @shadow I don't remember which but one of you said that but there was a rule about inadequate responses. Would like you to clarify to them about that. Jhonny can tag i never said he can't. just don't "spam" \(\color{#0cbb34}{\text{End of Quote}}\) He asked for her opinion she was just confirming that it's right, which was her opinion, that it was right i dont understand what your problem is with that? \(\color{#0cbb34}{\text{End of Quote}}\) exactly, but the opinion has to be more than i agree, like if thats what you are going to type then don't send, there has to be something that "contributes" @supie, this is a rule I saw either voca or dude post I am following that,

Vocaloid:

If you were specifically asked for your opinion on a post, that's fine. What we don't want is randoms barging in and saying "I agree" for SS boost. I'll double check w/ the other mods to be sure.

darkknight:

coolie

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight https://questioncove.com/users/dude#/updates/5fc7bc0b2f1d5c0a674eba00 \(\color{#0cbb34}{\text{End of Quote}}\) "If you were specifically asked for your opinion on a post, that's fine." Voca's words^

Vocaloid:

That being said, please do not abuse this function. Please only tag people if you genuinely need a second opinion.

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid If you were specifically asked for your opinion on a post, that's fine. What we don't want is randoms barging in and saying "I agree" for SS boost. I'll double check w/ the other mods to be sure. \(\color{#0cbb34}{\text{End of Quote}}\) ty. @Vocaloid but i think you know me and remember when i tag somebody i ask about what i answered or/and - to teach math - i like please/make they to contribute in how can be solved one posted math problem step by step easy understandably - why i tag much time @supie and @Laylalyssa so bc. they i think can be in future somebody who love math and can /wan teach math understandably step by step with/in collaboration with students who asked these posted problems - so than you think that this is a bad way how i thought - sorry - i dont make it in future but what and how made @darkknight above in this style - i dont believe that is a right way for the future of QC. - your opinion please ? ty

SemiDefinite:

the perfect storm

SemiDefinite:

OK, to solve this QU is to type the numbers into a basic ClassWiz and find you get a line thru' the O. All that argie for what 🤷‍♂️